OpenStudy (kamille):

Hello! I need to find\[\log _{5}6\] when\[\lg2=a\] and \[\lg3=b\]

OpenStudy (zehanz):

Assuming "lg" is also used for log, and its base is also 5, you can use the rule: \[\log_{b}p+\log_{b}q=\log_{b}pq\]By setting a=5, p=2 and q=3.

OpenStudy (kamille):

well, I have always thought that \[lga=\log _{10}a\] But thanks, anyway!

OpenStudy (zehanz):

You could be right. In that case, \(log_{5}6=\dfrac{lg6}{lg5}=\dfrac{lg2+lg3}{lg5}=\dfrac{a+b}{lg5}\). I used the rule \(log_{b}a=\dfrac{log_{c}a}{log_{c}b}\), to change the base from 5 to 10.

OpenStudy (kamille):

oh,thank you a lot!

OpenStudy (kamille):

the answer is \[\frac{ a+b }{ 1-a }\] maybe do you know how they get a-1?

OpenStudy (kamille):

oh, *1-a, my bad!

OpenStudy (kamille):

@mathslover any ideas?:/

mathslover (mathslover):

wait!

mathslover (mathslover):

btw lg means "log _ 2 "

mathslover (mathslover):

I searched net about it

mathslover (mathslover):

oh sorry \(lg = \log_{10} \)

OpenStudy (kamille):

well, I don know what lg means, we learnt about it at school. But i dont know how to get 1-a :o

OpenStudy (kamille):

I do know*

mathslover (mathslover):

wait I am getting it

OpenStudy (kamille):

okayy!

mathslover (mathslover):

what is log 10 (base 5) ?

OpenStudy (kamille):

10^x=5?

mathslover (mathslover):

see what I did : \[\large{\log_{10} 2 + \log_{10} 3 = \log_{10}(6)}\] \[\large{a + b = \log _{10}(6)}\] Changing base : \[\large{\log_{10}(6) = \frac{\log_5(6)}{\log_5{10}}}\] \[\large{(a+b)(\log_5(10))= \log_5(6)}\]

mathslover (mathslover):

gtg now, sorry I will discuss it later

OpenStudy (phi):

\[ \log_{5}(6)= \log_{5}(2 \cdot 3)=\log_{5}(2)+\log_{5}(3)\] now use the rule \[ \log_{a}(x)= \log_{a}(b)\log_{b}(x) \] with a=5, b=10 and x=3 we get \[ \log_{5}(3)= \log_{5}(10)\log_{10}(3) \] write 10 as 2*10 so we get \[ \log_{5}(3)= (\log_{5}(2)+ \log_{5}(5))\log_{10}(3) \] use the given \[ \log_{10}(3)=b \] and \[ \log_{5}(5)=1 \] to get \[ \log_{5}(3)= b\log_{5}(2)+ b \] now let's look at the other term \[ \log_{5}(2) = \log_{5}(10)\log_{10}(2) \] \[ \log_{5}(2) = (\log_{5}(2)+ \log_{5}(5))\log_{10}(2) \] use the given \[ \log_{10}(2)=a \] \[ \log_{5}(2) = a\log_{5}(2)+ a\] solve for \( \log_{5}(2) \) \[ \log_{5}(2)= \frac{a}{1-a} \] we now know enough to get the final answer

OpenStudy (zehanz):

Picking it up again from where I left: \(\log_{5}6=\dfrac{a+b}{\lg 5}\) (still assuming lg is log base 10). \(\log_{5}6=\dfrac{a+b}{\lg 5+\lg2-\lg2}\) \(\log_{5}6=\dfrac{a+b}{\lg(5 \cdot2)-\lg2}\) \(\log_{5}6=\dfrac{a+b}{\lg 10-\lg2}\) So finally we get: \(\log_{5}6=\dfrac{a+b}{1-a}\)

OpenStudy (phi):

\[ \log_{5}(2)+\log_{5}(3)= \frac{a}{1-a} +\frac{ab}{1-a}+ b \] which simplifies to your answer

OpenStudy (kamille):

thank you all!!